OpenStudy (anonymous):

Find the range of the function. f(x) = (x - 2)^2 + 2

OpenStudy (anonymous):

since \((x-2)^2\geq 0\) for all \(x\) (since it is a square) you have \[(x-2)^2+2\geq 2\]

OpenStudy (anonymous):

range is \(y\geq 0\) or \([2,\infty)\)